OpenStudy (tennisthegreat123):

10. Select the inequality that models the problem. The length of a rectangle is twice its width. The area of the rectangle is less than 18 square inches.a) 2x(2) < 18 b) x + 2x < 18 c) x(2x) > 18 In A, the 2 in parentheses is actually an exponent

OpenStudy (tennisthegreat123):

im thinking b?

OpenStudy (arthur_ser):

the length of is twice its width so L = 2W (L = length W = width) The area of a rectangle is L*W, in this case : area = W*2W=2W^2

OpenStudy (tennisthegreat123):

so it would be A/?

OpenStudy (arthur_ser):

yep

OpenStudy (tennisthegreat123):

tyvm